11.1The number of miles of useful tread wear (in 1000s) was determined for tires of each of five different makes of subcompact ca
Answer #1

Part (a)

i) Test statistic

= SSA/SSE

= 30.9/59.2

= 0.52 Answer 1

ii) p-value

= P(F4, 12 > 0.5220)

= 0.7216 First Option Answer 2

iii) Since p-value is very high (greater than any conceivable significance level), the null hypothesis of no difference in effect is accepted.

Hence Third Option Answer 3

Part (b)

Test statistic

= SSB/SSE

= 44.3/59.2

= 0.75 Answer 4

DONE

[Going beyond, (b) Part’s natural continuation

p-value

= P(F3, 12 > 0.7483)

= 0.5439

Conclusion: Since p-value is very high (greater than any conceivable significance level), the null hypothesis of no difference in effect is accepted.

Complete